OpenStudy (anonymous):

Is the set of multiples of 5 closed under division? Explain why or provide a counterexample if not. A. Yes, because the quotient of any two multiples of 5 is also a multiple of 5. B. Yes, because the quotient of any two multiples of 5 is also a multiple of 10. C. No, and a counterexample is 100 ÷ 5 = 20. D. No, and a counterexample is 25 ÷ 50 = . @misty1212

OpenStudy (anonymous):

and I can only choose one

OpenStudy (misty1212):

is it closed under division? if you take a multiple of 5, and divide it by another multiple of 5, must the result be a multiple of 5? hint, what is \(20\div5\)?

OpenStudy (anonymous):

20/5=4

OpenStudy (misty1212):

is 4 a multiple of 5?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

at least I don't think so

OpenStudy (misty1212):

right of course not

OpenStudy (anonymous):

so it is not a?

OpenStudy (misty1212):

so multiples of 5 are NOT closed under division \[20\div5=4\] is a "counterexample"

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

definitely not A or B

OpenStudy (anonymous):

oh and the answer to D. is 1/2

OpenStudy (misty1212):

yes it is

OpenStudy (anonymous):

so is the answer to the question D. or C.?

OpenStudy (misty1212):

D

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

\[25\div50=\frac{1}{2}\] is a "counter example"

OpenStudy (misty1212):

since \(\frac{1}{2}\) is not a multiple of 5

OpenStudy (anonymous):

can you explain why?

OpenStudy (misty1212):

i say "all cats are gray" you show me a cat that is black and white that is a "counterexample" to my claim that all cats are gray, one cat that is not gray

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what I don't get is how 100/5=20 is not a counter example

OpenStudy (misty1212):

because, dear, 20 IS a multiple of five i say all cats are gray, you show me a gray cat, that is not a counterexample

OpenStudy (anonymous):

oh ok